Given the following elements of an arithmetic series, a1=3 and a5=-17, find a21.
The correct answer is one of the following. Which one is correct?
A) -51
B) -100
C) -97
D) -92

Answer by greenestamps(13200)



Use the two given numbers in the sequence to find the common difference.

a5 is equal to a1 plus the common difference 5-1=4 times:

-17 = 3 + 4d
-20 = 4d
d = -5

Use that common difference and one of the given numbers to find the 21st term.

a21 is equal to a1 plus the common difference 21-1=20 times:

a21 = 3+20(-5) = 3-100 = -97